# FlowPath: Learning Data-Driven Manifolds with Invertible Flows for Robust Irregularly-sampled Time Series Classification

### v1

@@ Appendix A Theoretical Considerations
Thus $\mathcal{F}_{\Theta}$ is $L$-Lipschitz in $\theta$. Now, we exactly follow the standard argument in  (bartlett_spectrally-normalized_2017; chen_generalization_2020). If $\Theta$ has diameter $D$, covering‐number arguments yield
@@ Appendix B Implementation details of FlowPath
One strategy for robust NDEs is to constrain the learned dynamics through physical or mathematical priors on the vector field $f$. For example, stability can be enforced via conservation laws (white_stabilized_2023), or by designing stable classes of stochastic systems to mitigate noise sensitivity (oh_stable_2024), in the perspective of Neural Stochastic Differential Equations (Neural SDEs). Compared to that, FlowPath leaves the dynamics function $f$ unconstrained and instead achieves robustness by imposing a strong structural prior on the control path $\Phi(t)$ itself. This regularizes the system by ensuring the data manifold that drives the dynamics is well-behaved, rather than by limiting the flexibility of the dynamics that evolve upon it.
Another active area focuses on learning the data manifold explicitly, moving beyond the Euclidean assumptions of standard NDEs. One approach constrains the dynamics to remain tangent to a learned manifold, as in Neural Manifold ODEs (lou_neural_2020). Related advances in generative modeling apply continuous flows or diffusion processes on Riemannian manifolds (mathieu_riemannian_2020; huang_riemannian_2022), and recent work in metric flow matching emphasizes the importance of respecting intrinsic geometry (chen_flow_2024; kapusniak_metric_2024). FlowPath shares this manifold-aware motivation but with a more modular design. Our method learns a data-driven manifold to define the control path, while the dynamics themselves evolve in the ambient space of the NDE. This decouples the learning of the geometric structure from the evolution rule.
These different philosophies highlight a trade-off between expressivity and structure. Models like Characteristic-Neural ODEs (xu_characteristic_2023) are designed for maximum flexibility, enabling them to learn complex, non-homeomorphic transformations that can represent intersecting trajectories. In contrast, FlowPath intentionally leverages the structural constraint of a diffeomorphism as a powerful inductive bias. We posit that for robust classification from sparse observations, preserving the topology of the input data via an invertible map is a more beneficial constraint than allowing for arbitrary topological changes. This reflects a clear modeling choice, prioritizing a stable and well-behaved latent representation as the foundation for robust downstream performance.
@@
In our framework, the trainable control path $\Phi(t)$ is realized through an invertible mapping $F\colon[0,T]\times\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{x}}$ following bilos_neural_2021. Building on classical results in dynamical systems, we ensure $\Phi$ maintains bijectivity and differentiability, preserving local volumes in the latent space. We implement three architecture choices:
@@
Extending residual networks to continuous time (chen_neural_2018), we define:
@@
where $\varphi$ encodes time dependence and $f$ is a Lipschitz neural network. Invertibility follows from spectral normalization (behrmann_invertible_2019).
@@
Following brouwer_gru-ode-bayes_2019, we adapt the GRU architecture to continuous time:
@@
Inspired by normalizing flows (dinh_nice_2015; dinh_density_2017), we partition $\bm{z}$ and apply:
@@
We design our flow architecture to balance expressiveness with computational efficiency, guided by papamakarios_normalizing_2021’s normalizing-flow framework. In practice, we leverage Hutchinson’s trace estimator (hutchinson_stochastic_1989) for efficient Jacobian-related computations, further reducing the overhead of flow-based modeling (grathwohl_ffjord_2019; bilos_neural_2021; oh_dualdynamics_2025). The flow architecture is designed to balance computational efficiency and representational expressiveness. We optimize its configuration during the tuning process and validate our choices through an ablation study.
@@
The adjoint method is a powerful technique often used to calculate the gradients of systems governed by differential equations (pollini_adjoint_2018). The adjoint-sensitive method, proposed to solve Neural ODEs, reduces the memory requirements and computational burden that would otherwise be associated with backpropagating gradients through the solution of an ODE or CDE (chen_neural_2018; kidger_neural_2020). The primary concept of the adjoint-sensitive method is to introduce an adjoint state, usually denoted as $\bm{a}(t)$, which satisfies an auxiliary differential equation running backward in time. Algorithm 1 summarizes the procedure for performing the parameter update using the adjoint-sensitivity framework in FlowPath.
@@ Appendix C Detailed results on time series classification with missingness
We performed classification experiments using 18 datasets spanning three distinct domains from the University of East Anglia (UEA) and the University of California Riverside (UCR) Time Series Classification Repository [^1] http://www.timeseriesclassification.com/ (bagnall_uea_2018; h_a_dau_ucr_2019). These experiments utilized the Python library sktime (loning_sktime_2019). As summarized in Table 8, datasets in the Sensor domain contain only univariate time series, while datasets in the other domains include both univariate and multivariate time series. To address the issue of varying sequence lengths, all time series were uniformly scaled to match the length of the longest sequence (oh_stable_2024; oh_dualdynamics_2025). This scaling approach ensured consistent input dimensions across datasets for modeling.
After aligning the series lengths, random missing values, including 30%, 50%, and 70%, were introduced into each time series to simulate real-world scenarios. The datasets were then partitioned into training, validation, and testing subsets, using a ratio of 70%, 15%, and 15% respectively. Unlike the original splits from bagnall_uea_2018, which varied across datasets, this partitioning ensured a uniform evaluation framework. Finally, the altered time series for each variable were integrated into a unified dataset. To evaluate model robustness, multiple cross-validation iterations were performed, each with a unique random seed for variability and statistical reliability.
@@
In every model and dataset, we seek to minimize the validation loss by optimizing the hyperparameters as suggested by oh_stable_2024. The hyperparameters are optimized using the following scheme: Learning rate $\eta\in\{10^{-4},10^{-1}\}$, sampled from a log-uniform distribution; Number of layers $n_{l}\in\{1,2,3,4\}$, determined via grid search; Hidden vector dimensions $n_{h}\in\{16,32,64,128\}$, selected through grid search; Batch size chosen from $\{16,32,64,128\}$, adjusted according to the total data size.
It is important that for the sake of fairness and integrity in comparative analysis, we employed the original source code provided by the authors of the aforementioned models. In regard to hyperparameter tuning, we undertook a systematic approach for each model and dataset. This was facilitated by the employment of the Python library ray [^2] https://github.com/ray-project/ray (moritz_ray_2018; liaw_tune_2018). The hyperparameters were calibrated based on the original dataset, and the same configuration was consistently applied to scenarios incorporating missing data, ensuring consistency in experimental settings.
For RNN-based methods, these hyperparameters are used in the fully-connected layer, while for NDE-based methods, they are employed in the embedding layer and vector fields. Furthermore, Euler(-Maruyama) method is used to solve NDE-based methods. In case of the Neural Flow and the proposed FlowPath, three different flow models are considered with original source code  [^3] https://github.com/mbilos/neural-flows-experiments. As suggested in bilos_neural_2021, we optimize the best flow model for each dataset. Also, we compare the performance of the different flow models in the ablation study. Each model undergoes training for a total of 100 epochs, with the best-performing model selected based on the lowest validation loss. The early-stopping is applied when the loss does not improve for 10 consecutive epochs.
@@
We perform pairwise comparisons between FlowPath and each baseline model across four missingness levels, as well as an aggregated setting that includes all levels. Statistical tests use one-sided Wilcoxon signed-rank tests (demsar_statistical_2006), with Holm-Bonferroni adjustment for multiple comparisons (giacalone_bonferroni-holm_2018). Significance is determined at the adjusted threshold of $p<0.05$ (benavoli_should_2016).
Table 13: Pairwise comparison of FlowPath against the all baselines across missing rates using Wilcoxon signed-rank tests
@@ Appendix D Detailed results on real-world dataset with sensor dropout
The PAMAP2 dataset (a_reiss_introducing_2012) [^4] https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring contains daily activity data from nine individuals, recorded using three inertial measurement units (placed on the wrist, chest, and ankle) and a heart rate monitor. To evaluate model robustness under sensor dropout scenarios, we adapted this dataset following the preprocessing method proposed by zhang_graph-guided_2022555https://github.com/mims-harvard/Raindrop.
@@
Results of hyperparameter tuning.
Following zhang_graph-guided_2022, training was performed with a batch size of 128, a learning rate of 0.0001, and 20 epochs. Similar to the previous experiment, we employed grid search for hyperparameter tuning: number of layers $n_{l}\in\{1,2,3,4\}$, and hidden vector dimensions $n_{h}\in\{16,32,64,128\}$. The results of this tuning are presented in Table 14, where we discovered that a configuration with $n_{l}=4$ layers and $n_{h}=128$ hidden vector dimensions yields the optimal performance of F1 Score.
Settings Neural CDE FlowPath
Spline MLP ResNet GRU Coupling
Regular 94.2 $\pm$0.5 93.3 $\pm$0.3 94.1 $\pm$0.4 94.8 $\pm$0.2 94.4 $\pm$0.1
(a) Accuracy
Settings Neural CDE FlowPath
Spline MLP ResNet GRU Coupling
Regular 95.2 $\pm$0.4 94.2 $\pm$0.2 94.8 $\pm$0.4 95.8 $\pm$0.4 95.1 $\pm$0.3
(b) Precision
Settings Neural CDE FlowPath
Spline MLP ResNet GRU Coupling
Regular 94.8 $\pm$0.5 93.6 $\pm$0.6 94.4 $\pm$0.7 95.5 $\pm$0.2 94.6 $\pm$0.3
(c) Recall
Settings Neural CDE FlowPath
Spline MLP ResNet GRU Coupling
Regular 95.0 $\pm$0.4 93.9 $\pm$0.4 94.6 $\pm$0.5 95.6 $\pm$0.2 94.8 $\pm$0.2
(d) F1 Score

### v2

@@ Appendix A Theoretical Considerations
Thus $\mathcal{F}_{\Theta}$ is $L$-Lipschitz in $\theta$. Now, we exactly follow the standard argument in  (Bartlett et al. 2017; Chen et al. 2020). If $\Theta$ has diameter $D$, covering‐number arguments yield
@@ Appendix B Implementation details of FlowPath
One strategy for robust NDEs is to constrain the learned dynamics through physical or mathematical priors on the vector field $f$. For example, stability can be enforced via conservation laws (White et al. 2023), or by designing stable classes of stochastic systems to mitigate noise sensitivity (Oh et al. 2024), in the perspective of Neural Stochastic Differential Equations (Neural SDEs). Compared to that, FlowPath leaves the dynamics function $f$ unconstrained and instead achieves robustness by imposing a strong structural prior on the control path $\Phi(t)$ itself. This regularizes the system by ensuring the data manifold that drives the dynamics is well-behaved, rather than by limiting the flexibility of the dynamics that evolve upon it.
Another active area focuses on learning the data manifold explicitly, moving beyond the Euclidean assumptions of standard NDEs. One approach constrains the dynamics to remain tangent to a learned manifold, as in Neural Manifold ODEs (Lou et al. 2020). Related advances in generative modeling apply continuous flows or diffusion processes on Riemannian manifolds (Mathieu and Nickel 2020; Huang et al. 2022), and recent work in metric flow matching emphasizes the importance of respecting intrinsic geometry (Chen and Lipman 2024; Kapuśniak et al. 2024). FlowPath shares this manifold-aware motivation but with a more modular design. Our method learns a data-driven manifold to define the control path, while the dynamics themselves evolve in the ambient space of the NDE. This decouples the learning of the geometric structure from the evolution rule.
These different philosophies highlight a trade-off between expressivity and structure. Models like Characteristic-Neural ODEs (Xu et al. 2023) are designed for maximum flexibility, enabling them to learn complex, non-homeomorphic transformations that can represent intersecting trajectories. In contrast, FlowPath intentionally leverages the structural constraint of a diffeomorphism as a powerful inductive bias. We posit that for robust classification from sparse observations, preserving the topology of the input data via an invertible map is a more beneficial constraint than allowing for arbitrary topological changes. This reflects a clear modeling choice, prioritizing a stable and well-behaved latent representation as the foundation for robust downstream performance.
@@
In our framework, the trainable control path $\Phi(t)$ is realized through an invertible mapping $F\colon[0,T]\times\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{x}}$ following Bilos et al. (2021). Building on classical results in dynamical systems, we ensure $\Phi$ maintains bijectivity and differentiability, preserving local volumes in the latent space. We implement three architecture choices:
@@
Extending residual networks to continuous time (Chen et al. 2018), we define:
@@
where $\varphi$ encodes time dependence and $f$ is a Lipschitz neural network. Invertibility follows from spectral normalization (Behrmann et al. 2019).
@@
Following Brouwer et al. (2019), we adapt the GRU architecture to continuous time:
@@
Inspired by normalizing flows (Dinh et al. 2015, 2017), we partition $\bm{z}$ and apply:
@@
We design our flow architecture to balance expressiveness with computational efficiency, guided by Papamakarios et al. (2021)’s normalizing-flow framework. In practice, we leverage Hutchinson’s trace estimator (Hutchinson 1989) for efficient Jacobian-related computations, further reducing the overhead of flow-based modeling (Grathwohl et al. 2019; Bilos et al. 2021; Oh et al. 2025b). The flow architecture is designed to balance computational efficiency and representational expressiveness. We optimize its configuration during the tuning process and validate our choices through an ablation study.
@@
The adjoint method is a powerful technique often used to calculate the gradients of systems governed by differential equations (Pollini et al. 2018). The adjoint-sensitive method, proposed to solve Neural ODEs, reduces the memory requirements and computational burden that would otherwise be associated with backpropagating gradients through the solution of an ODE or CDE (Chen et al. 2018; Kidger et al. 2020). The primary concept of the adjoint-sensitive method is to introduce an adjoint state, usually denoted as $\bm{a}(t)$, which satisfies an auxiliary differential equation running backward in time. Algorithm 1 summarizes the procedure for performing the parameter update using the adjoint-sensitivity framework in FlowPath.
@@ Appendix C Detailed results on time series classification with missingness
We performed classification experiments using 18 datasets spanning three distinct domains from the University of East Anglia (UEA) and the University of California Riverside (UCR) Time Series Classification Repository [^1] http://www.timeseriesclassification.com/ (Bagnall et al. 2018; H. A. Dau et al. 2019). These experiments utilized the Python library sktime (Löning et al. 2019). As summarized in Table 8, datasets in the Sensor domain contain only univariate time series, while datasets in the other domains include both univariate and multivariate time series. To address the issue of varying sequence lengths, all time series were uniformly scaled to match the length of the longest sequence (Oh et al. 2024, 2025b, 2025c). This scaling approach ensured consistent input dimensions across datasets for modeling.
After aligning the series lengths, random missing values, including 30%, 50%, and 70%, were introduced into each time series to simulate real-world scenarios. The datasets were then partitioned into training, validation, and testing subsets, using a ratio of 70%, 15%, and 15% respectively. Unlike the original splits from Bagnall et al. (2018), which varied across datasets, this partitioning ensured a uniform evaluation framework. Finally, the altered time series for each variable were integrated into a unified dataset. To evaluate model robustness, multiple cross-validation iterations were performed, each with a unique random seed for variability and statistical reliability.
@@
In every model and dataset, we seek to minimize the validation loss by optimizing the hyperparameters as suggested by Oh et al. (2024). The hyperparameters are optimized using the following scheme: Learning rate $\eta\in\{10^{-4},10^{-1}\}$, sampled from a log-uniform distribution; Number of layers $n_{l}\in\{1,2,3,4\}$, determined via grid search; Hidden vector dimensions $n_{h}\in\{16,32,64,128\}$, selected through grid search; Batch size chosen from $\{16,32,64,128\}$, adjusted according to the total data size.
It is important that for the sake of fairness and integrity in comparative analysis, we employed the original source code provided by the authors of the aforementioned models. In regard to hyperparameter tuning, we undertook a systematic approach for each model and dataset. This was facilitated by the employment of the Python library ray [^2] https://github.com/ray-project/ray (Moritz et al. 2018; Liaw et al. 2018). The hyperparameters were calibrated based on the original dataset, and the same configuration was consistently applied to scenarios incorporating missing data, ensuring consistency in experimental settings.
For RNN-based methods, these hyperparameters are used in the fully-connected layer, while for NDE-based methods, they are employed in the embedding layer and vector fields. Furthermore, Euler or Euler–Maruyama method is used to solve NDE-based methods. In case of the Neural Flow and the proposed FlowPath, three different flow models are considered with original source code  [^3] https://github.com/mbilos/neural-flows-experiments. As suggested in Bilos et al. (2021), we optimize the best flow model for each dataset. Also, we compare the performance of the different flow models in the ablation study. Each model undergoes training for a total of 100 epochs, with the best-performing model selected based on the lowest validation loss. The early-stopping is applied when the loss does not improve for 10 consecutive epochs.
@@
We perform pairwise comparisons between FlowPath and each baseline model across four missingness levels, as well as an aggregated setting that includes all levels. Statistical tests use one-sided Wilcoxon signed-rank tests (Demšar 2006), with Holm-Bonferroni adjustment for multiple comparisons (Giacalone et al. 2018). Significance is determined at the adjusted threshold of $p<0.05$ (Benavoli et al. 2016).
@@
Table 13: Pairwise comparison of FlowPath against all baselines across missing rates using Wilcoxon signed-rank tests
@@ Appendix D Detailed results on real-world dataset with sensor dropout
The PAMAP2 dataset (A. Reiss and D. Stricker 2012) [^4] https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring contains daily activity data from nine individuals, recorded using three inertial measurement units (placed on the wrist, chest, and ankle) and a heart rate monitor. To evaluate model robustness under sensor dropout scenarios, we adapted this dataset following the preprocessing method proposed by Zhang et al. (2022) [^5] https://github.com/mims-harvard/Raindrop.
@@
(a) Accuracy
(b) Precision
(c) Recall
(d) F1 Score
Results of hyperparameter tuning.
Following Zhang et al. (2022), training was performed with a batch size of 128, a learning rate of 0.0001, and 20 epochs. Similar to the previous experiment, we employed grid search for hyperparameter tuning: number of layers $n_{l}\in\{1,2,3,4\}$, and hidden vector dimensions $n_{h}\in\{16,32,64,128\}$. The results of this tuning are presented in Table 14, where we discovered that a configuration with $n_{l}=4$ layers and $n_{h}=128$ hidden vector dimensions yields the optimal performance of F1 Score.
Settings Neural CDE FlowPath
Spline MLP ResNet GRU Coupling
Regular 94.2 $\pm$0.5 93.3 $\pm$0.3 94.1 $\pm$0.4 94.8 $\pm$0.2 94.4 $\pm$0.1
(a) Accuracy
Settings Neural CDE FlowPath
Spline MLP ResNet GRU Coupling
Regular 95.2 $\pm$0.4 94.2 $\pm$0.2 94.8 $\pm$0.4 95.8 $\pm$0.4 95.1 $\pm$0.3
(b) Precision
Settings Neural CDE FlowPath
Spline MLP ResNet GRU Coupling
Regular 94.8 $\pm$0.5 93.6 $\pm$0.6 94.4 $\pm$0.7 95.5 $\pm$0.2 94.6 $\pm$0.3
(c) Recall
Settings Neural CDE FlowPath
Spline MLP ResNet GRU Coupling
Regular 95.0 $\pm$0.4 93.9 $\pm$0.4 94.6 $\pm$0.5 95.6 $\pm$0.2 94.8 $\pm$0.2
(d) F1 Score